myininaya (myininaya):

This one looks weird: \[f(x)=\lim_{n \rightarrow \infty} \frac{x^{2n}-1}{x^{2n}+1} \\ \text{ where is } f \text{ continuous ?}\]

OpenStudy (anonymous):

i wanna cry ='(

myininaya (myininaya):

lol I'm thinking about it too...

myininaya (myininaya):

If this helps or if this is a hint to anyone, this comes after sequences and series discussion in the james stewart 6e calculus book.

OpenStudy (anonymous):

i felt series involved hmm

OpenStudy (anonymous):

ok can i switch the limit to 0 ?

myininaya (myininaya):

i guess if we do a substitution

OpenStudy (anonymous):

hehe

myininaya (myininaya):

lol but you wanted to say f(x)=(1-1)/(1+1)=0

myininaya (myininaya):

not sure if this helps but we can write as \[f(x)=\lim_{n \rightarrow \infty}\frac{1-\frac{1}{x^{2n}}}{1+\frac{1}{x^{2n}}}\]

OpenStudy (anonymous):

\(\lim_{n \to 0}\frac{x^{\frac{1}{2n}}-1}{x^{\frac{1}{2n}}+1} \)

OpenStudy (anonymous):

no we cant do what u posted :O

OpenStudy (anonymous):

can we ? :O

OpenStudy (anonymous):

ok my brain is out of service literally , ill wait until u reveal some solution :)

myininaya (myininaya):

I wonder if we should just consider some random numbers... \[x=-1 \\ f(-1)=\lim_{n \rightarrow \infty}\frac{(-1)^n-1}{(-1)^n+1}\] so this limit wouldn't exist so it isn't defined at x=-1

myininaya (myininaya):

but f(-1/2)=-1

OpenStudy (anonymous):

interesting ,,,,

myininaya (myininaya):

i cheated

myininaya (myininaya):

so f(x)=-1 for when -1<x<1

myininaya (myininaya):

i don't know how to show that though yet and also i guess wolfram considered all cases when finding limits and the other limits didn't exist

OpenStudy (anonymous):

oh :O

OpenStudy (anonymous):

ok then i'll work on this ltr

myininaya (myininaya):

if n goes to infinity then (x^2n) goes to 0 if -1<x<1

ganeshie8 (ganeshie8):

\[\lim \limits_{n\to \infty} x^n = \infty , ~0\]

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}x^{2n} =0 \\ \text{ for } x \in (-1,1) \\ f(x)=\frac{0-1}{0+1}=-1\] for x=-1 to x=1

ganeshie8 (ganeshie8):

does that mean the function is continuous everywhere except -1 and 1 ? idk, a real analyst will have a more rigorous argument for proving continuity @eliassaab @ikram002p @zzr0ck3r

OpenStudy (anonymous):

\(\frac{x^{2n}-1}{x^{2n}+1} = 1-\frac{2}{x^{2n}+1} \)

OpenStudy (anonymous):

im here -.-

OpenStudy (anonymous):

\(\Large \frac{x^{2n}-1}{x^{2n}+1} = 1-\frac{2}{x^{2n}+1} \) ok this should help

myininaya (myininaya):

and if we consider anything outside..maybe we can try to use l'hospital I guess and see what we come up with \so x>1... \[\lim_{n \rightarrow \infty} \frac{x^n \ln(x)-0}{x^n \ln(x)+0}=1 , x>0\] wait this isn't right is it?

OpenStudy (anonymous):

\( \lim 1-\frac{2}{x^{2n}+1} =1-2=-1 \) done yollaaa

myininaya (myininaya):

i mean x>1 *

myininaya (myininaya):

oops forgot the 2

OpenStudy (anonymous):

@ganeshie8 no its cont at 1,-1

myininaya (myininaya):

\[\lim_{n \rightarrow \infty} \frac{2x^{2n} \ln(x)-0}{2x^{2n} \ln(x)+0}=1 , x>1 \]

OpenStudy (anonymous):

im here :(

OpenStudy (anonymous):

that sounds cool @myininaya i forgot that its a limit

ganeshie8 (ganeshie8):

Oh so it is continuous everywhere ?

myininaya (myininaya):

so I wonder if it is continuous on (-1,1) U (1,inf)

myininaya (myininaya):

well we can't do that if x<-1

myininaya (myininaya):

because of the ln part

myininaya (myininaya):

so we would have to consider something else for x<-1 I guess

OpenStudy (anonymous):

x^2n>0 ppl so its cont !

OpenStudy (anonymous):

2n even power :O

OpenStudy (anonymous):

im i missing something ?

OpenStudy (anonymous):

so what its cont :O whats that cool program xD

myininaya (myininaya):

so the function was continuous everywhere i feel like i have been tricked

ganeshie8 (ganeshie8):

i stil think it can get discont at x = -1 and 1, not sure hmm

ganeshie8 (ganeshie8):

we need to consider n->infinity

OpenStudy (anonymous):

its even power xD

OpenStudy (anonymous):

If you divide top and bottom by \(x^{-2n}\), It's just a matter of when that is not continuous, I think.

OpenStudy (anonymous):

check demo. only

OpenStudy (anonymous):

Since \(1^{\infty}\) is an indeterminate, form, I am guessing it has issues at \(x=\pm 1\).

OpenStudy (anonymous):

ok i'll tell u something it would be continues and smooth but derivative does not exist , happy now ?

ganeshie8 (ganeshie8):

show me analysis argument marki

OpenStudy (anonymous):

like this 

OpenStudy (anonymous):

im too tired

OpenStudy (anonymous):

it has contour track :'(

ganeshie8 (ganeshie8):

it is just "drawn" continuously without lifting the hand but everyone knows a square wave has jump discontinuity at transitions

OpenStudy (anonymous):

with no branch cut

myininaya (myininaya):

stupid question how do you find f'(x)?

myininaya (myininaya):

like can you ignore the limit part for the time being and just use quotient rule?

OpenStudy (anonymous):

\[ \lim_{x\to 1^+}f(x) = 1 \neq -1 = \lim_{x\to 1^{-}}f(x) \]

myininaya (myininaya):

and then reinsert limit part

ganeshie8 (ganeshie8):

wio did u look at the animation ? how can u say that without fixing n value ?

myininaya (myininaya):

\[f(x)=\lim_{n \rightarrow a}g(x,n) \\ \text{ \in general what is the rule here for finding derivative } \]

myininaya (myininaya):

I'm not sure why I want to find the derivative lol

OpenStudy (anonymous):

Well \(x\to 1^+\) would mean that we treat limit as if \(x>1\) and \(x^{-2n}\to 0\): \[ \lim_{n\to \infty }\frac{x^{2n}-1}{x^{2n}+1} = \lim_{n\to \infty }\frac{1-x^{-2n}}{1+x^{-2n}} =1 \]

OpenStudy (anonymous):

Is that not good enough?

myininaya (myininaya):

and so since f(1)=0 then it is not continuous at x=1

OpenStudy (anonymous):

I don't think you can say \(f(1)=0\).

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @myininaya \[f(x)=\lim_{n \rightarrow a}g(x,n) \\ \text{ \in general what is the rule here for finding derivative } \] \(\color{blue}{\text{End of Quote}}\) In this case, you differentiate with respect to \(n\).

OpenStudy (anonymous):

the sun is rising here finally XD

myininaya (myininaya):

\[f(1)=\lim_{n \rightarrow \infty}\frac{(1)^{2n}-1}{(1)^{2n}+1}=\lim_{n \rightarrow \infty}\frac{1-1}{1+1}=\frac{0}{2}=0\] why can't we do this?

OpenStudy (anonymous):

Oh, I guess you can do that.

OpenStudy (anonymous):

yes yes '!

myininaya (myininaya):

and you showed \[\lim_{x \rightarrow 1^+}f(x)=1 \neq f(1)\]

myininaya (myininaya):

so this was a double limit question

OpenStudy (anonymous):

my problem is with the smoothly of this function , its so smooth xD

OpenStudy (anonymous):

like x=1 function

myininaya (myininaya):

well in @ganeshie8 picture when the n's were getting larger the turned at x=1 and x=-1 were getting sharper

myininaya (myininaya):

the turns*

OpenStudy (anonymous):

somehow when u convert smooth dot derivable functions to complex they also continues

myininaya (myininaya):

so for large n it is not smooth there at x=1 or x=-1

OpenStudy (anonymous):

lets try this definition of continuity  no branches cuts ...